LIST OF PARTICIPANTS

Abreu, M. (IST, Lisboa, Portugal) mabreu@math.ist.pt

Albuquerque, R. P. R. (Univ. Évora, Portugal) ribeiro@ptmat.lmc.fc.ul.pt

Alias, L. J. (Univ. Murcia, Spain) ljalias@fcu.es

Armario, J. A. (Esc. Univ. Ing. T. Agrícola, Sevilla, Spain) armario@cica.es

Ballmann, W. (Univ. Bonn, Germany) ballmann@math.uni-bonn.de

Barreiro, M. E. (Univ. Coimbra, Portugal) mefb@mat.uc.pt

Bonome, A. (Univ. Santiago, Spain) bonome@zmat.usc.es

Breda, A. M. (Univ. Aveiro, Portugal) ambreda@mat.ua.pt

Caetano, A. J. (Univ. Lisboa, Portugal) pcaetano@lmc.fc.ul.pt

Carter, S. (Univ. Leeds, U.K.) s.carter@leeds.ac.uk

Carvalho, M. A. (Univ. Algarve, Portugal) mcarva@mozart.si.ualg.pt

Caseiro, R. (Univ. Coimbra, Portugal) raquel@mat.uc.pt

Castro, I. (Univ. Jaén, Spain) icastro@ujaen.es

Chaves, R. (Univ. São Paulo, Brazil) rosab@ime.usb.br Cordero, L. A. (Univ. Santiago, Spain) cordero@zmat.usc.es

Craveiro de Carvalho, F. (Univ. Coimbra, Portugal) fjcc@mat.uc.pt

Da Silva, I. (Univ. Lisboa, Portugal) isilva@alfi.cii.fc.ul.pt

De Andres, L. (Univ. País Vasco, Bilbao, Spain) mtpandol@lg.ehu.es

Dodson, C. T. J. (Univ. Manchester, U. K.) dodson@umist.ac.uk

Eells, J. (Univ. Cambridge, U. K.) je208@cus.cam.ac.uk

Farran, H. (Univ. Kuwait, Kuwait) Farran@MATH-1.sci.kuniv.edu

Ferreira, M. C. (Univ. Lisboa, Portugal) cecilia@ptmat.lmc.fc.ul.pt

Ferreira, M. J. (Univ. Lisboa, Portugal) mjferr@ptmat.fc.ul.pt

Galvão, M. E. (Univ. São Paulo, Brazil) elisa@ime.usb.br

Garcia-rio, E. (Univ. Santiago, Spain) eduardo@zmat.usc.es

Gonzalez Davilla, J. (Univ. La Laguna, Tenerife, Spain) jcgonzalez@ull.es

Gothen, P. (Univ. Porto, Portugal) pbgothen@fc .up.pt

Guedes de Oliveira, A. (Univ. Porto, Portugal) agoliv@fc.up.pt

Hervella, L. (Univ. Santiago, Spain) hervella@zmat.usc.es Ibanez, R. (Univ. País Vasco, Bilbao, Spain) mtbibtor@lg.ehu.es

Leandro, C. (Univ. Minho, Portugal) leandro@math.uminho.pt

Liberman, P. (Univ. Paris VII, France)

Machado, A. (Univ. Lisboa, Portugal) armac@lmc.fc.ul.pt

Macho-Stadler, M. (Univ. País Vasco, Bilbao, Spain) mtpmastm@lg.ehu.es

Magalhães, L. (Univ. Porto, Portugal) mlsmagal@fc.up.pt

Martin, F. (Univ. Granada, Spain) fmartin@goliat.ugr.es

Matsushita, Y. (Univ. Shiga, Japan) yasuo@mech.usp.ac-jp

Medrano, S. L. (Univ. Mexico, Mexico) santiago@matem.unam.mx

Mena, F. (Univ. Minho, Portugal) fmena@math.uminho.pt

Moreira, L. (Univ. Porto, Portugal) mlmoreir@fc.up.pt

Nelli, B. (Univ. Pisa, Italy)

Nunes da Costa, J. M. (Univ. Coimbra, Portugal) jmcosta@mat.uc.pt

Oliveira, J. M. (Univ. Minho, Portugal) jmo@math.uminho.pt

Osserman, R. (Univ. Berkeley, U.S.A.) osserman@msri.org

Palmer, W. B. (Univ. Durham, U. K.)

Patrício, J. P. (Univ. Minho, Portugal) pedro@math.uminho.pt

Pereira do Vale, A. (Univ. Minho, Portugal) avale@math.uminho.pt

Perez, J. (Univ. Granada, Spain) jperez@goliat.ugr.es

Perez Lopez, T. (Univ. Santiago, Spain)

Pinto, M. R. (Univ. Porto, Portugal) mspinto@fc.up.pt

Robertson, S. A. (Univ. Southampton, U. K.) Sar.Rob@btinternet.com

Rocha, J. (Univ. Porto, Portugal) jrocha@fc.up.pt

Rodriguez, J. L. (Univ. Barcelona, Spain) jlrodri@mat.uab.es

Romero-Fuster, M. C. (Univ. Valencia, Spain) Carmen.Romero@post.uv.es

Salamon, D. (Univ. Warwick, U. K.) das@maths.warwick.ac.uk

Salavessa, I. (I. S. T., Lisboa, Portugal) isabel@cartan.ist.utl.pt

Sanabria, E. (Univ. Valencia, Spain) sanabria@uv.es

Sanchez, I. (Univ. Granada, Spain) ignacios@goliat.ugr.es

Smith., P. (Univ. Minho, Portugal) psmith@math.uminho.pt

Tarrio, A. (Univ. Coruna, Spain) madorana@udc.es Teles Correia, J. M. (Univ. Coimbra, Portugal) jteles@mat.uc.pt

Ugarte, L. (Univ. Zaragoza, Spain) mtpugvil@lg.ehu.es

Vasquez-Aval, M. E. (Univ. Santiago, Spain) meva@zmat.usc.es

Vazquez, L. R. (Univ. Santiago, Spain) rvazquez@abonados.cplus.es

Verjosky, A. (Univ. Lille, France) Alberto.Verjovsky@univ-lille1.fr

Wegner, B. (T. Univ. Berlin, Germany) wegner@math.tu-berlin.de

West, A. (Univ. Leeds, U. K.) a.west@leeds.ac.uk